OpenStudy (anonymous):

Help Please! 0.25 moles of hydrogen carbon x with a molar mass of 72 m/mol was burnt completely in an excess of oxygen.The volume of CO2 formed was 30 dm3. i) What is the molar formula of the hydrocarbon x? 2a)Cyclo hexane has the molecular formula C6H10 whereas cyclo hexene has the molecular formula C6H12.Compare these formulae with those of a straight chain alkanes and alkenes containing 6.comment on thne differences found. iii)State and explain How you would distinguish between two hydro carbonsstated in 2a) above

OpenStudy (anonymous):

'72m /mol' can u state what is the unit 'm' mean?

OpenStudy (anonymous):

Sorry that should be 'g'

OpenStudy (anonymous):

so '72g/mol'

OpenStudy (anonymous):

\[30dm ^{3}/24dm ^{3} = 1.25moles\] mass of carbon from the carbon dioxide formed is : \[1.25moles \times 6g= 7.5 g\] for the hydrocarbon, 0.25 moles and 72g as molar mass, we can find its relative mass: \[0.25moles \times 72g/moles = 18g\] which mean in 18 gram of hydrocarbon, the carbon dioxide produced contains 7.5 g of carbon, while mass of H in the hydrocarbon is 10.5g. (noted that 10.5+7.5=18) but the answer turn out a bit awkward, 7.5g/6g = moles of carbon = 1.25 moles while 10.5g/1g =10.5 moles of H. the molar formula is \[C_{5}H_{42} ??\] not sure about it

OpenStudy (anonymous):

Thank you so much.